40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Jul 29 22:08:32 2010, Alberto, Update, SUS, More optics installed on the BS table 
    Reply  Fri Jul 30 02:54:04 2010, Koji, Update, 40m Upgrading, More optics installed on the BS table 
Message ID: 3329     Entry time: Fri Jul 30 02:54:04 2010     In reply to: 3326
Author: Koji 
Type: Update 
Category: 40m Upgrading 
Subject: More optics installed on the BS table 

July 29 Thu [Steve, Alberto, Kiwamu, Koji]

We placed some optics in the BS chamber.
The chambers are ready to be pumped down on Friday once the heavy door is placed.

- Clean room work

  • Engraved two Y2 mirrors and PBS@532nm
  • Engraved three DLC mounts
  • Each of the mounts needs a 3.5 inch post. We found there is no stock of the post in the lab! Also the clamps!
  • Took the posts from the temporarily removed optics although we need to return those optics into the table during the next vent.
  • We should count the # of the mounts and count the needed posts. Posts and clamps can be either a DLC thick post or New Focus pedestal.

- In the chamber

  • The terminal holder was moved as Alberto described
  • The green steering optics were placed as Alberto described
  • Note: the PBS is flipped in the mount (reflection side is back side)
  • The table leveling
  • Releasing EQ stops / Check OSEMs / Adjust OSEMs (BS OSEMs are untouched)

- After closing the chamber

  • The BS OSEM mumbo-jumbo

Quote:

[Koji, Steve, Kiwamu, Alberto]

- This afternoon we installed a few new optics on the BS table: GR_PBS, GRY_SM2, GRY_SM1.

- We pulled up the cables so that we had more freedom to move one of the cable towers farther South.

- Then we re-leveled the table. PRM OSEMs were adjusted to be nominal insertions.

- Koji released the earthquake stops on BS but the readout of the OSEMs was apparently frozen on the MEDM screens.
Initially we thought it was a software problem. a nuclear reboot didn't solve it. We spent the following three hours investigating the cause.
Eventually it turned out that the earthquake stops on BS weren't actually fully released.

We opened the tank and accessed to BS. Releasing the earthquake stops in full solved the issue. The OSEMs readout went back to normal.

 

ELOG V3.1.3-